# Neural networks for the approximation of Euler’s elastica

### v1

@@ 5.1 Numerical experiments with $q_{\boldsymbol{\rho}}^{\textrm{c}}$
As for the case of the discrete network, we perform an in-depth investigation of this learning setting by varying the architecture of the continuous neural network and the hyperparameters in the training procedure, whose range of options can be found in Table 6. In this case, we define the loss as in Equation (12), with $\gamma=10^{-2}$. The weight decay is systematically set to 00.
Hyperparameter
Range
Distribution
architecture
{MLP, ResNet, MULT}
discrete uniform
normalisation
{True, False}
discrete uniform
activation function $\sigma$
{Tanh, Swish, Sigmoid, Sine}
discrete uniform
#layers $\ell$
$\{3,\ldots,8\}$
discrete uniform
#hidden nodes in each layer
$[10,200]\cap\mathbb{N}$
discrete uniform
learning rate $\eta$
$[1\cdot 10^{-4},1\cdot 10^{-1}]$
log uniform
Table 6: Hyperparameter ranges for the continuous network $q_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set with $90\%\--10\%$ splitting into training and test set. The first column of the table reports the hyperparameters and network architectures we test for. The second describes the set of allowed values for each, while the third specifies how such values are explored through Optuna. The weight decay is systematically set to 00. MULT stands for multiplicative neural network and corresponds to the network in Equations (20)-(24) of Appendix A.
Table 7 collects the combination of hyperparameters yielding the best results on the both-ends data set. This leads to a training error equal to $1.869\cdot 10^{-6}$ and a test error equal to $4.81\cdot 10^{-6}$. In Figure 3, the comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$ is shown. As we can see in the plot showing the mean error over the trajectories, the error on the end nodes is nonzero, since we are not imposing boundary conditions by construction. This is in contrast to the corresponding plot for the discrete network in Figure 1.
Selected hyperparameters
architecture
MULT
normalisation
True
activation function $\sigma$
Tanh
#layers $\ell$
5
#hidden nodes in each layer
190
learning rate $\eta$
$3.025\cdot 10^{-3}$
Table 7: Combination of hyperparameters yielding the best results for the continuous network $q_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set with $90\%\--10\%$ splitting into training and test set. The results are shown in Figure 3.
Figure 3: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$ for the continuous network $q_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set with $90\%\--10\%$ splitting into training and test set. These results are obtained with the hyperparameters from Table 7, that yield a training error equal to $1.869\cdot 10^{-6}$ and a test error equal to $4.810\cdot 10^{-6}$. For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
@@
Data set splittingTraining - $\begin{array}[]{c}\text{Data set splitting}\\ \text{Training \-- test}\end{array}$
Training accuracy Test accuracy
10% - 10%
$2.383\cdot 10^{-4}$
$7.784\cdot 10^{-4}$
20% - 10%
$5.612\cdot 10^{-5}$
$7.285\cdot 10^{-5}$
40% - 10%
$7.104\cdot 10^{-6}$
$9.275\cdot 10^{-6}$
90% - 10%
$1.869\cdot 10^{-6}$
$4.810\cdot 10^{-6}$
Table 8: Behaviour of the continuous network $q_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set with fewer training data points. The size of the training set varies, while that of the test set is fixed. The last row corresponds to the results in Figure 3.
@@ 5.2 Numerical experiments with $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$
Here we consider a neural network approximation of the angle $\theta(s)$ that parametrises the tangent vector $\mathbf{q}^{\prime}(s)=(\cos(\theta(s)),\sin(\theta(s)))$. By design, the approximation $\tau_{\boldsymbol{\rho}}^{\mathrm{c}}$ of the tangent vector $\mathbf{q}^{\prime}$ satisfies the constraint $\|\tau_{\boldsymbol{\rho}}^{\mathrm{c}}(\boldsymbol{x}^{i})(s)\|_{2}=1$ for every $s\in[0,L]$ and $\boldsymbol{x}^{i}\in\mathbb{R}^{8}$. We also analyse how the neural network approximation behaves when the boundary conditions $\tau_{\boldsymbol{\rho}}^{\mathrm{c}}(\boldsymbol{x}^{i})(0)=\mathbf{q}^{\prime}(0)$ and $\tau_{\boldsymbol{\rho}}^{\mathrm{c}}(\boldsymbol{x}^{i})(L)=\mathbf{q}^{\prime}(L)$ are imposed by construction. To do so, we model the parametric function $\hat{\theta}_{\boldsymbol{\rho}}^{\textrm{c}}$, defined in Equation (13), in one of the two following ways:
$\hat{\theta}_{\boldsymbol{\rho}}^{\textrm{c}}(\boldsymbol{x}^{i})(s)=f_{\boldsymbol{\rho}}(s,\theta_{0}^{i},\theta_{N}^{i}),$ (17)
$\begin{split}\hat{\theta}_{\boldsymbol{\rho}}^{\textrm{c}}(\boldsymbol{x}^{i})(s)&=f_{\boldsymbol{\rho}}(s,\theta_{0}^{i},\theta_{N}^{i})+(\theta_{0}^{i}-f_{\boldsymbol{\rho}}(0,\theta_{0}^{i},\theta_{N}^{i}))e^{-100s^{2}}\\ &+(\theta_{N}^{i}-f_{\boldsymbol{\rho}}(L,\theta_{0}^{i},\theta_{N}^{i}))e^{-100(s-L)^{2}},\end{split}$ (18)
@@
where $f_{\boldsymbol{\rho}}:\mathbb{R}^{3}\to\mathbb{R}$ is any neural network, and we recall that $\pi(\boldsymbol{x}^{i})=(\theta_{0}^{i},\theta_{N}^{i})$. We remark that, in the case of the parameterisation in Equation (18), one gets $\theta_{\boldsymbol{\rho}}^{\textrm{c}}(\boldsymbol{x}^{i})(0)=\theta_{0}^{i}$ and $\theta_{\boldsymbol{\rho}}^{\textrm{c}}(\boldsymbol{x}^{i})(L)=\theta_{N}^{i}$ up to machine precision, due to the fast decay of the Gaussian function. As in the previous sections, we collect the hyperparameter and architecture options with the respective range of choices in Table 9, and we report the results without imposing the boundary conditions in Figure 4, while those imposing them in Figure 5, in both cases using the both-ends data set, with $90\%\--10\%$ splitting into training and test set. The results shown in the two figures correspond respectively to training errors of $5.821\cdot 10^{-6}$ and $6.068\cdot 10^{-6}$, and test errors of $6.231\cdot 10^{-6}$ and $6.289\cdot 10^{-6}$. The best performing hyperparameter combinations can be found in Tables 10 and 11.
Hyperparameter
Range
Distribution
architecture
{MLP, ResNet, MULT}
discrete uniform
activation function $\sigma$
{Tanh, Swish, Sigmoid}
discrete uniform
#layers $\ell$
$\{3,\ldots,8\}$
discrete uniform
#hidden nodes in each layer
$[50,200]\cap\mathbb{N}$
discrete uniform
learning rate $\eta$
$[1\cdot 10^{-3},1\cdot 10^{-1}]$
log uniform
Table 9: Hyperparameter ranges for the continuous network $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set. The first column of the table reports the hyperparameters and network architectures we test for. The second describes the set of allowed values for each, while the third specifies how such values are explored through Optuna. The weight decay is systematically set to 00, and no normalisation is applied.
@@
Figure 4: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$, for the case $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ is modelled as in Equation (17), with $90\%\--10\%$ splitting of the both-ends data set into training and test set. These results are obtained with the hyperparameters from Table 10, that yield a training error equal to $5.821\cdot 10^{-6}$ and a test error equal to $6.231\cdot 10^{-6}$. For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
Selected hyperparameters
architecture
MULT
activation function $\sigma$
Tanh
#layers $\ell$
$7$
#hidden nodes in each layer
$143$
learning rate $\eta$
$3.007\cdot 10^{-3}$
Table 10: Combination of hyperparameters yielding the best results for the case when $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ is modelled as in Equation (17), with $90\%\--10\%$ splitting of the both-ends data set into training and test set. The results are shown in Figure 4.
@@
Figure 5: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$, for the case $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ is modelled as in Equation (18), with $90\%\--10\%$ splitting of the both-ends data set into training and test set. These results are obtained with the hyperparameters from Table 11, that yield a training error equal to $6.068\cdot 10^{-6}$ and a test error equal to $6.289\cdot 10^{-6}$. For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
Selected hyperparameters
architecture
MULT
activation function $\sigma$
Tanh
#layers $\ell$
$4$
#hidden nodes in each layer
$175$
learning rate $\eta$
$1.846\cdot 10^{-3}$
Table 11: Combination of hyperparameters yielding the best results for the case when $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ is modelled as in Equation (18), with $90\%\--10\%$ splitting of the both-ends data set into training and test set. The results are shown in Figure 5.
@@ 6 Discussion
The results in Figures 4 and 5 are comparable, especially looking at the mean error plots. This suggests that the imposition of the boundary conditions, in the proposed way, is not limiting the expressivity of the considered network. Thus, given the boundary value nature of our problem, these figures advocate the enforcement of the boundary conditions on the network $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$. However, due to the chosen reconstruction procedure in Equation (15) for the variable $\mathbf{q}$, we are able to impose the boundary conditions on $\mathbf{q}$ only on the left node. Clearly, other more symmetric reconstruction procedures can be adopted, but the proposed one has proved to provide better experimental results.
Comparing the results related to $q_{\boldsymbol{\rho}}^{\textrm{c}}$ with those of $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$, we notice similar performances in terms of training and test errors. In both the cases, they have one order of magnitude more than the corresponding training and test errors of the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$. Thus, as a results of our experiments, we can conclude that
@@
The total accuracy error of a neural network model can be defined by splitting it into three components: approximation error, optimisation error, and generalisation error (see, e.g., [17]). To achieve excellent agreement between predicted and reference trajectories, it is important to select the appropriate architecture and fine-tune the model hyperparameters. Our results demonstrate that we can construct a network that is expressive enough to provide a small approximation error and with very good generalisation capability.
@@ 6.1 Future work
In the methods presented in this paper, there is no interaction between the mathematical problem and the neural network model once the data set is created. As a way to improve the results presented here, one could include the Euler elastica model directly into the training process. This could be done either by directly imposing in the loss function that $\mathbf{q}(s)$ satisfies the differential equations (5), or one could add the constrained action integral from Equation (4) into the loss function that is minimised, see e.g. [28, 10, 12, 11].
There are many promising directions in order to follow up this work. One is to consider 3D versions of the Euler elastica, another is to look at the dynamical problem, and finally one may examine industrial applications. As an example, we refer to the modelling of endoscopes due to the high bending deformation of these medical devices under certain loading cases [30]. The approximation of the elastica through neural networks can indeed help in the prediction of the deformed configuration of the beam in constrained narrow environments.
@@
Appendix A Other neural network architectures
Another example of neural network architecture, besides the MLP defined in Equation (9), is the residual neural network (ResNet), where
$f_{j}^{RES}(\mathbf{x})=\mathbf{x}+\mathbf{B}_{j}^{T}\sigma(\mathbf{A}_{j}\mathbf{x}+\mathbf{b}_{j})\in\mathbb{R}^{n},$ (19)
with $\mathbf{x}\in\mathbb{R}^{n}$, $\mathbf{A}_{j},\mathbf{B}_{j}\in\mathbb{R}^{n_{j}\times n}$, $\mathbf{b}_{j}\in\mathbb{R}^{n_{j}}$, and $\boldsymbol{\rho}=\{\mathbf{A}_{j},\mathbf{b}_{j},\mathbf{B}_{j}\}_{j=1}^{\ell}$.
We also provide the expression of the forward propagation of the multiplicative network used for the experiments in Section 5:
$\mathbf{U}=\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1}),\quad\mathbf{V}=\phi(\mathbf{W}_{2}\mathbf{x}+\mathbf{b}_{2})$ (20)
@@ Appendix A Other neural network architectures
$f_{\boldsymbol{\rho}}^{MULT}(\mathbf{x})=\mathbf{W}\mathbf{H}_{\ell+1}+\mathbf{b},$ (24)
@@
Appendix B Further results on the hyperparameters of the neural networks
$\begin{array}[]{c}\text{Hyperparameter}\\ \text{combination}\end{array}$ Data set splitting
10% - 10%
20% - 10%
40% - 10%
90% - 10%
architecture
MLP
MLP
MLP
MLP
normalization
False
False
False
True
activation function $\sigma$
Tanh
Tanh
Tanh
Tanh
number of layers $\ell$
4
4
3
4
#hidden nodes in each layer
950
351
904
879
learning rate $\eta$
$1.019\cdot 10^{-3}$
$5.455\cdot 10^{-3}$
$1.429\cdot 10^{-3}$
$1.378\cdot 10^{-3}$
weight decay
$1.79\cdot 10^{-6}$
$2.142\cdot 10^{-7}$
$1.317\cdot 10^{-7}$
$1.535\cdot 10^{-7}$
$\gamma$
$8.595\cdot 10^{-3}$
$2.973\cdot 10^{-3}$
$9.338\cdot 10^{-3}$
$4.242\cdot 10^{-3}$
batch size
64
32
32
32
Table 12: Choice of hyperparameters for the training of the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends data set with different numbers of training data points.
$\begin{array}[]{c}\text{Hyperparameter}\\ \text{combination}\end{array}$ Data set splitting
10% - 10%
20% - 10%
40% - 10%
90% - 10%
architecture
MULT
MULT
MULT
MULT
normalization
True
True
True
True
activation function $\sigma$
Tanh
Tanh
Sine
Tanh
number of layers $\ell$
5
6
6
5
#hidden nodes in each layer
193
121
169
190
learning rate $\eta$
$4.548\cdot 10^{-3}$
$4.913\cdot 10^{-3}$
$4.2\cdot 10^{-3}$
$3.025\cdot 10^{-3}$
Table 13: Choice of hyperparameters for the training of the continuous network $q_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set with different numbers of training data points. The weight decay is systematically set to 0.

### v2

@@ 5.1 Numerical experiments with $q_{\boldsymbol{\rho}}^{\textrm{c}}$
As for the case of the discrete network, we perform an in-depth investigation of this learning setting. In this case, the experiments are run on a GPU-P100 machine. For this continuous setup, the standard MLP architecture does not provide accurate results even after a hyperparameter optimisation routine. Given, hence, that the simple MLP architecture does not seem to be flexible enough to capture the complexity of the elastica solution in this continuous framework, we move to a different architecture that we call MULT for the presence of multiplicative interactions in its architecture. This network has demonstrated superior performance to standard fully connected neural networks in the context of operator learning, see e.g. [51]. Details on this architecture can be found in Appendix A. We fix the learning rate $\eta$ to $5\cdot 10^{-3}$ and only vary the number of layers and of hidden nodes in the training procedure, with the range of options reported in Appendix B, Table 6. In this case, we define the loss as in Equation (12), with $\gamma=10^{-2}$. The weight decay is systematically set to 00. For the both-ends data set, this leads to a training error equal to $3.554\cdot 10^{-6}$, a validation error equal to $4.779\cdot 10^{-6}$, and a test error equal to $4.354\cdot 10^{-6}$. In Figure 3, the comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$ is shown. As we can see in the plot showing the mean error over the trajectories, the error on the end nodes is nonzero, since we are not imposing boundary conditions by construction. This is in contrast to the corresponding plot for the discrete network in Figure 1.
Figure 3: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$ for the continuous network $q_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set with $80\%\--10\%\--10\%$ splitting into training, validation, and test sets. The mean squared error on the test set equals $4.354\cdot 10^{-6}$. For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
@@
Data set splittingTraining - validation - $\begin{array}[]{c}\text{Data set splitting}\\ \text{Training \-- validation \-- test}\end{array}$ Training accuracy Validation accuracy Test accuracy
10% - 10% - 10% $2.146\cdot 10^{-4}$ $1.252\cdot 10^{-3}$ $8.811\cdot 10^{-4}$
20% - 10% - 10% $4.187\cdot 10^{-5}$ $4.239\cdot 10^{-5}$ $6.279\cdot 10^{-5}$
40% - 10% - 10% $7.037\cdot 10^{-6}$ $8.357\cdot 10^{-6}$ $8.434\cdot 10^{-6}$
80% - 10% - 10% $3.554\cdot 10^{-6}$ $4.779\cdot 10^{-6}$ $4.354\cdot 10^{-6}$
Table 3: Behaviour of the continuous network $q_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set with fewer training data points. The size of the training set varies, while that of the validation and the test sets is fixed. The last row corresponds to the results in Figure 3.
@@ 5.2 Numerical experiments with $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$
Here we consider a neural network approximation of the angle $\theta(s)$ that parametrises the tangent vector $\mathbf{q}^{\prime}(s)=(\cos(\theta(s)),\sin(\theta(s)))$. By design, the approximation $\tau_{\boldsymbol{\rho}}^{\mathrm{c}}$ of the tangent vector $\mathbf{q}^{\prime}$ satisfies the constraint $\|\tau_{\boldsymbol{\rho}}^{\mathrm{c}}(\mathbf{x}^{i})(s)\|_{2}=1$ for every $s\in[0,L]$ and $\mathbf{x}^{i}\in\mathbb{R}^{8}$. We also analyse how the neural network approximation behaves when the boundary conditions $\tau_{\boldsymbol{\rho}}^{\mathrm{c}}(\mathbf{x}^{i})(0)=\mathbf{q}^{\prime}(0)$ and $\tau_{\boldsymbol{\rho}}^{\mathrm{c}}(\mathbf{x}^{i})(L)=\mathbf{q}^{\prime}(L)$ are imposed by construction. To do so, we model the parametric function $\hat{\theta}_{\boldsymbol{\rho}}^{\textrm{c}}$, defined in Equation (13), in one of the two following ways:
$\hat{\theta}_{\boldsymbol{\rho}}^{\textrm{c}}(\mathbf{x}^{i})(s)=f_{\boldsymbol{\rho}}(s,\theta_{0}^{i},\theta_{N}^{i}),$ (17)
$\begin{split}\hat{\theta}_{\boldsymbol{\rho}}^{\textrm{c}}(\mathbf{x}^{i})(s)&=f_{\boldsymbol{\rho}}(s,\theta_{0}^{i},\theta_{N}^{i})+(\theta_{0}^{i}-f_{\boldsymbol{\rho}}(0,\theta_{0}^{i},\theta_{N}^{i}))e^{-100s^{2}}\\ &+(\theta_{N}^{i}-f_{\boldsymbol{\rho}}(L,\theta_{0}^{i},\theta_{N}^{i}))e^{-100(s-L)^{2}},\end{split}$ (18)
@@
where $f_{\boldsymbol{\rho}}:\mathbb{R}^{3}\to\mathbb{R}$ is any neural network, and we recall that $\pi(\mathbf{x}^{i})=(\theta_{0}^{i},\theta_{N}^{i})$. We remark that, in the case of the parameterisation in Equation (18), one gets $\theta_{\boldsymbol{\rho}}^{\textrm{c}}(\mathbf{x}^{i})(0)=\theta_{0}^{i}$ and $\theta_{\boldsymbol{\rho}}^{\textrm{c}}(\mathbf{x}^{i})(L)=\theta_{N}^{i}$ up to machine precision, due to the fast decay of the Gaussian function. As in the previous sections, we collect the hyperparameter and architecture options with the respective range of choices in Table 8, and we report the results without imposing the boundary conditions in Figure 4, while those imposing them in Figure 5, in both cases using the both-ends data set, with $80\%\--10\%\--10\%$ splitting into training, validation, and test sets. The results shown in the two figures correspond respectively to training errors of $6.288\cdot 10^{-6}$ and $5.301\cdot 10^{-6}$, validation errors $5.874\cdot 10^{-6}$ and $5.065\cdot 10^{-6}$, and test errors of $5.089\cdot 10^{-6}$ and $4.385\cdot 10^{-6}$. The best-performing hyperparameter combinations can be found in Table 9.
@@
Figure 4: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$, for the case $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ is modelled as in Equation (17), with $80\%\--10\%\--10\%$ splitting of the both-ends data set into training, validation, and test sets. The mean squared error on the test set equals $6.289\cdot 10^{-6}$. For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
@@
Figure 5: Comparison over test trajectories for $\mathbf{q}$ and $\mathbf{q}^{\prime}$, for the case $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ is modelled as in Equation (18), with $80\%\--10\%\--10\%$ splitting of the both-ends data set into training, validation, and test sets. The mean squared error on the test set equals $4.385\cdot 10^{-6}$. For presentation purposes, only 10 randomly selected trajectories are considered in the first two plots.
@@ 6 Discussion
The results in Figures 4 and 5 are comparable, especially looking at the mean error plots. This suggests that the imposition of the boundary conditions, in the proposed way, is not limiting the expressivity of the considered network. Thus, given the boundary value nature of our problem, these figures advocate the enforcement of the boundary conditions on the network $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$. However, due to the chosen reconstruction procedure in Equation (15) for the variable $\mathbf{q}$, we are able to impose the boundary conditions on $\mathbf{q}$ only on the left node. Other more symmetric reconstruction procedures can be adopted, but the proposed one has provided better experimental results.
Comparing the results related to $q_{\boldsymbol{\rho}}^{\textrm{c}}$ with those of $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$, we notice similar performances in terms of training and test errors. In both cases, they have one order of magnitude more than the corresponding training and test errors of the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$. Thus, as a result of our experiments, we can conclude that
@@
The total accuracy error of a neural network model can be defined by splitting it into three components: approximation error, optimisation error, and generalisation error (see e.g. [17]). To achieve excellent agreement between predicted and reference trajectories, it is crucial to select the appropriate architecture and fine-tune the model hyperparameters. Our results demonstrate that we can construct a network that is expressive enough to provide a small approximation error and with very good generalisation capability.
Lastly, we have compared the time cost of the Neural Network prediction against the traditional approach with numerical solvers as described in Section 2.2. The discrete and the continuous approaches outperformed the traditional solvers with an average speedup of 105.000 times and 260.000 times, respectively, across the test trajectories. The training time of the continuous network is 1.25 times larger than that of the discrete network. It’s important to note that these results are subject to certain limitations, such as the specific choice of the hyperparameters or the machine used to train and test the network. These findings suggest that using neural networks to predict new solutions of the elastica for unseen boundary conditions is much more time efficient than the classical numerical methods, although requiring intensive offline training.
@@ 6.1 Future work
In the methods presented in this paper, the mathematical problem and the neural network model do not interact once the data set is created. To improve the results presented here, one could include Euler’s elastica model directly into the training process. This could be done either by directly imposing in the loss function that $\mathbf{q}(s)$ satisfies the differential equations (5), or one could add the constrained action integral from Equation (4) into the loss function that is minimised, see e.g. [30, 10, 12, 11].
There are many promising directions to follow up on this work. One is to consider 3D versions of Euler’s elastica, another is to look at the dynamical problem, and finally one may examine industrial applications, as mentioned in the introduction.
CRediT author statement.
Elena Celledoni: Conceptualisation, Validation, Writing - Review & Editing, Supervision, Funding acquisition. Ergys Çokaj: Validation, Investigation, Visualisation, Writing - Review & Editing. Andrea Leone: Methodology, Software, Investigation, Writing - Original Draft, Review & Editing. Sigrid Leyendecker: Conceptualisation, Writing - Review & Editing, Supervision, Funding acquisition. Davide Murari: Methodology, Software, Investigation, Writing - Original Draft, Review & Editing. Brynjulf Owren: Conceptualisation, Validation, Writing - Review & Editing, Supervision, Funding acquisition. Rodrigo T. Sato Martín de Almagro: Methodology, Validation, Writing - Review & Editing, Supervision. Martina Stavole: Methodology, Software, Investigation, Writing - Original Draft, Review & Editing.
@@
Appendix A Architecture for the continuous network
We provide the expression of the forward propagation of the multiplicative network MULT used for the experiments in Section 5:
$\mathbf{U}=\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1}),\quad\mathbf{V}=\sigma(\mathbf{W}_{2}\mathbf{x}+\mathbf{b}_{2})$ (19)
@@ Appendix A Architecture for the continuous network
$f_{\boldsymbol{\rho}}^{\mathrm{MULT}}(\mathbf{x})=\mathbf{W}\mathbf{H}_{\ell+1}+\mathbf{b},$ (23)
@@
Appendix B Details on hyperparameter optimisation
We provide here further details on the hyperparameter optimisation strategy with Optuna, relative to the results in Sections 4 and 5. The tables below display the hyperparameters we optimise for in each of the networks, the ranges and the distribution, as well as the selected combinations used to perform the experiments in the paper.
Hyperparameter Range Distribution
#layers $\ell$$\{0,...,10\}$ discrete uniform
#hidden nodes $[10,1000]\cap\mathbb{N}$ discrete uniform
$\gamma$ $[0,1\cdot 10^{-2}]$ uniform
Table 4: Hyperparameter ranges for the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$. The first column of the table reports the hyperparameters we test. The second describes the set of allowed values for each, while the third specifies how such values are explored through Optuna.
$\begin{array}[]{c}\text{Hyperparameter}\\ \text{combination}\end{array}$ of trajectories of the whole dataset in the training set
10% 20% 40% 80%
# layers $\ell$ 4 4 4 4
#hidden nodes 950 978 997 985
$\gamma$ $7.044\cdot 10^{-3}$ $6.336\cdot 10^{-3}$ $9.004\cdot 10^{-3}$ $3.853\cdot 10^{-3}$
Table 5: Choice of hyperparameters for the training of the discrete network $q_{\boldsymbol{\rho}}^{\textrm{d}}$ tested on the both-ends data set with different sizes of the training set, with the validation and test sets each containing $10\%$ of trajectories of the dataset.
Hyperparameter Range Distribution
#layers $\ell$$\{5,\ldots,10\}$ discrete uniform
#hidden nodes $[10,250]\cap\mathbb{N}$ discrete uniform
Table 6: Hyperparameter ranges for the continuous network $q_{\boldsymbol{\rho}}^{\textrm{c}}$. The first column of the table reports the hyperparameters we test. The second describes the set of allowed values for each, while the third specifies how such values are explored through Optuna.
$\begin{array}[]{c}\text{Hyperparameter}\\ \text{combination}\end{array}$ of trajectories of the whole dataset in the training set
10% 20% 40% 80%
# layers $\ell$ 6 7 8 6
#hidden nodes 139 185 181 106
Table 7: Choice of hyperparameters for the training of the continuous network $q_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set with different sizes of the training set, with the validation and test sets each containing $10\%$ of trajectories of the dataset.
Hyperparameter Range Distribution
#layers $\ell$$\{1,\ldots,10\}$ discrete uniform
#hidden nodes $[50,200]\cap\mathbb{N}$ discrete uniform
Table 8: Hyperparameter ranges for the continuous network $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$.The first column of the table reports the hyperparameters we test. The second describes the set of allowed values for each, while the third specifies how such values are explored through Optuna.
$\begin{array}[]{c}\text{Hyperparameter}\\ \text{combination}\end{array}$ $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$
$\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ as in (17) $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ as in (18)
# layers $\ell$ 8 8
#hidden nodes 93 58
Table 9: Choice of the hyperparameters for the training of the continuous network $\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ tested on the both-ends data set with $80\%\--10\%\--10\%$ splitting into training, validation, and test sets. The second column shows the combination of hyperparameters yielding the best result corresponding to Figure 4, while the third column that corresponding to Figure 5.